Find a equation in slope intercept form for the line through (6,-2) with a slope of 1\2
Answer by lwsshak3(11628)

Find a equation in slope intercept form for the line through (6,-2) with a slope of 1\2
**
standard form of equation for a straight line: y=mx+b, m=slope, b= y-intercept
for given problem:
y=x/2+b
solving for b using coordinates of given point (6,-2)
-2=6/2+b
-2=3+b
b=-5
Equation:
y=x/2-5